3. Cross-rate Bid-Ask Quotes. National Bank quotes the following rates for the Euro and Yen:

EUR/NZD Bid: 0.5153 Ask: 0.5168

JPY/NZD Bid: 81.14 Ask: 81.32

Calculate the bid/ask quotes for JPY/EUR.
Answer #1

INZD - 0.5153 - 0.5168 Euro INZD = 81.14 - 81.32 JPY 1 JPY = ? Euro 1 JPY = Euro x NZD NZD JPY Bid = 0,5153 x L = 0.0063367 8